OpenStudy (anonymous):

what is the arc length when θ = 2pi/3 and the radius is 8 cm

OpenStudy (anonymous):

26pi/3 cm 10pi/3 cm 16pi/3 cm pi/3 cm

OpenStudy (anonymous):

Arc Length \[A = r \theta \]

OpenStudy (anonymous):

yeah i got 16.7 but it isnt an answer choice

OpenStudy (anonymous):

Make sure your calculator is in radians mode.

OpenStudy (anonymous):

Just plug R = 8 and theta = 2pi/3.

OpenStudy (anonymous):

Where Theta Is In Radians.

OpenStudy (anonymous):

You really don'\[\frac{ 8 }{ 1} * \frac{ 2\pi }{3 }\]t need a calculator for this

OpenStudy (anonymous):

thanks man